# Eulerian magnitude homology: diagonality, injective words, and regular path homology

### v1

@@ 5. The injective nerve of categories and the regular magnitude-path spectral sequence
The goal of this section is construct the regular magnitude-path spectral sequence, in analogy with the magnitude-path spectral sequence developed in [Asa23]. To do so, it shall be more convenient to use the categorical language, and to generalize the framework of eulerian magnitude homology to quivers. In the first subsection we introduce the tool of injective nerves of categories, and then we shall apply it to the reachable categories, getting the regular magnitude-path spectral sequence. As an application of the construction, we shall provide some computations of regular path homology of regularly diagonal graphs.
@@ 5.1. The injective nerve
Let $\mathbf{2}$ denote the category consisting of the objects $E$ and $V$, with two non-identity morphisms $s,t\colon E\to V$ called the source and the target. We shall also represent, and identify, a quiver $Q$ by the corresponding directed graph with set of vertices $V$, and set of edges $E$. Let $\mathbf{Fin}$ be the full subcategory of $\mathbf{Set}$ of finite sets. By a (finite) quiver we shall mean a functor $Q\colon\mathbf{2}\to\mathbf{Fin}$. Morphisms of quivers are natural transformations of functors. The category $\mathbf{Quiver}$ of finite quivers and morphisms of quivers is the functor category $\mathbf{Fun}(\mathbf{2},\mathbf{Fin})$.
@@
between the category of quivers and the 1-category $\mathbf{Cat}$ of small categories and functors. In this adjunction, the functor $\mathrm{Free}$ is the free functor that turns a quiver into a category, and $U$ is the corresponding forgetful functor.
Recall that an ordered $k$-clique of a directed graph $G$ is a totally ordered $k$-tuple $(v_{1},...,v_{k})$ of vertices of $G$ with the property that, for every $i<j$, the pair $(v_{i},v_{j})$ is an ordered edge of $G$. This definition extends to quivers in the natural way and, for any given quiver $Q$, we can construct the associated directed flag complex $\mathrm{dFl}(Q)$ as well. Furthermore, the directed flag complex yields a functor $\mathrm{dFl}\colon\mathbf{Quiver}\to\mathbf{OSC}$, where $\mathbf{OSC}$ is the category of ordered simplicial complexes.
Let $\Delta\mathbf{Set}$ be the category of semi-simplicial sets, and $\mathbf{sSet}$ the category of simplicial sets. Recall that a semi-simplicial set is a sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of sets together with functions called face maps between them which encode that an element in $X_{n+1}$ has $n+1$ faces (i.e. boundary segments), and that such faces are elements of $X_{n}$; similarly, a simplicial set is a collection of sets $\{S_{n}\}_{n\in\mathbb{N}}$ together with face maps $f_{i}\colon S_{i+1}\to S_{i}$ and degeneracy maps $s_{i}\colon S_{i}\to S_{i+1}$, where the $f_{i}$’s are constructed similarly to the semi-simplicial case, and the $s_{i}$’s encode which $(i+1)$-simplices are really just $n$-simplices regarded as degenerate $(n+1)$-simplices. Also, recall that an ordered simplicial complex yields a semi-simplicial set, and that each semi-simplicial set yields a simplicial set. We consider the following composition
[…]
of functors, where $\mathbf{OSC}\hookrightarrow\Delta\mathbf{Set}\hookrightarrow\mathbf{sSet}$ are the standard inclusions (see, eg[Rie14, Section 1.5] for a detailed account of adjunctions involving simplicial sets).
@@
For a category $\mathbf{C}$, $N^{\iota}(\mathbf{C})$ is the injective nerve of $\mathbf{C}$.
Taking injective nerves is functorial on the category $\mathbf{Cat}$ of categories and functors. We can characterize the simplices of the injective nerve as the functors that are injective on objects. First, consider the category $[n]$, which is the category with objects the natural numbers $0,\dots,n$ and with a morphism $i\to j$ if and only if $i\leq j$. Saying otherwise, $[n]$ is the free category generated by the transitive tournament $T_{n}$. Denote by $\mathbf{Cat}^{\iota}(\mathbf{D},\mathbf{C})$ the family of functors from $\mathbf{D}$ to $\mathbf{C}$ which are injective on objects.
@@
Recall the adjunction of Eq. (6). The set of ordered $(n+1)$-cliques of the quiver $U(\mathbf{C})$ is given by the ordered tuples $(c_{0},\dots,c_{n})$ consisting of objects of $\mathbf{C}$, with the further requirement that for all $i<j$ there is a morphism $f_{i,j}\colon c_{i}\to c_{j}$ in $\mathbf{C}$. Then, each such tuple defines a functor $S\colon[n]\to\mathbf{C}$ with $S(i)\coloneqq c_{i}$ and $S(i<j)\coloneqq f_{i,j}$. The functor $S$ is injective on objects because the set of vertices is ordered (and without repetitions). Vice versa, every functor $S\colon[n]\to\mathbf{C}$ which is injective on objects yields an ordered $(n+1)$-clique, hence an $n$-simplex of the directed flag complex – and, consequently, of $N^{\iota}(\mathbf{C})$. ∎
In view of Lemma 5.2, we have that the injective nerve is the simplicial set associated to the semi-simplicial set with simplices the tuples $(c_{0},\dots,c_{k})$ of distinct objects of $\mathbf{C}$.
We now shall justify the term injective nerve used in Definition 5.1. Let $\Delta$ be the simplex category, and denote by $\Delta_{+}$ the wide subcategory of $\Delta$ containing only the injective functions. Alternatively, $\Delta_{+}$ is (equivalent to) the category of finite totally ordered sets and order-preserving injections. Recall that for a given (small) category $\mathbf{C}$, the nerve $N(\mathbf{C})$ is the simplicial set
@@
where $\mathbf{Cat}(-,\mathbf{C})$ denotes the Hom-functor. If we restrict to the category $\Delta_{+}$, we get the nerve of $\mathbf{C}$ as a semi-simplicial set – which means, without taking degeneracies. We shall further restrict to the set of functors $\mathbf{Cat}^{\iota}(-,\mathbf{C})$ which are injective on objects; hence, we consider the composition
@@
The injective nerve $N^{\iota}(\mathbf{C})$ of a small category $\mathbf{C}$ is the simplicial set associated to the composition
@@
By Lemma 5.2, the set of non-degenerate simplices of the injective nerve of a category $\mathbf{C}$ can be described as the set of functors $[n]\to\mathbf{C}$ which are injective on objects. The $i$-th face map operator in $\Delta_{+}$ acts by forgetting the $i$-th entry in the tuples $(c_{0},\dots,c_{k})$. Hence, the face maps are coherent with the face maps of the directed flag complex construction. The statement follows after considering the associated simplicial sets. ∎
As a consequence, Proposition 5.3 can be taken as alternative definition of the simplicial nerve; we shall use both viewpoints.
Let $\mathbf{Poset}$ be the category of posets, and poset maps. Each poset can be seen as a category in the standard way, and there is an inclusion functor $\mathbf{Poset}\hookrightarrow\mathbf{Cat}$.
Proposition 5.4.
The diagram
${\mathbf{Poset}}$${\mathbf{Cat}}$${\mathbf{sSet}}$$\scriptstyle{N}$$\scriptstyle{N^{\iota}}$
is commutative.
Proof.
Let $P$ be a poset, seen as a category. The set of functors $[n]\to P$ which are injective on objects is the same set of all possible functors $[n]\to P$, since $P$ has no non-trivial loops. The statement follows from the definitions. ∎
@@
The nerve of a category is a fully faithful functor $\mathbf{Cat}\to\mathbf{sSet}$. As a consequence of Proposition 5.3, the injective nerve $N^{\iota}$ is a faithful functor, but this is generally not full – unless we restrict to the category of posets.
We may ask whether the injective nerve of a category is a (simplicial) nerve, at the cost of changing the base category. However, the following result tells us that in general the injective nerve is not the nerve of any category – nor of any quasicategory.
@@
Consider the category with two objects $0,1$, and non-trivial morphisms $0\to 1$, $1\to 0$. Consider the inner horn $\lambda^{n}_{i}$ in the case $n=2$, $i=1$ given by the arrows $0\to 1$, $1\to 0$. Then, $\lambda^{n}_{i}$ has no lift to $\Delta^{2}$ – because the simplex $(0,1,0)$ does not belong to $N^{\iota}(\mathbf{C})$. ∎
The inclusion $\mathbf{Cat}^{\iota}(-,\mathbf{C})\subseteq\mathbf{Cat}(-,\mathbf{C})$ induces a natural transformation
$N^{\iota}(\mathbf{C})\to N(\mathbf{C})$
which is an embedding of simplicial sets. As mentioned in the introduction, it would be interesting to know when the inclusion
$N^{\iota}(\mathbf{C})\to N(\mathbf{C})$
is a weak equivalence of simplicial sets.
@@ 5.2. The regular magnitude-path spectral sequence
We are interested in the injective nerve of reachable categories, whose definition we now recall (see [CR24]). Let $Q$ be a finite directed graph or quiver.
@@
$\mathbf{Reach}_{Q}(v,w)\coloneqq\begin{cases}*&\text{if there is a path from $v$ to $w$ in $Q$};\\ \emptyset&\text{ otherwise}.\end{cases}$
@@
Recall the adjunction between quivers and categories of Eq. (6), where $U\colon\mathbf{Cat}\to\mathbf{Quiver}$ denotes the forgetful functor.
@@
If $G$ is a directed tree, then $\mathrm{Free(G)}$ and $\mathbf{Reach}_{G}$ are isomorphic categories. More generally, they are isomorphic if and only if the quiver $Q$ does not contain directed cycles nor quasi-bigons – see [CR24, Proposition 3.8]. In such case, the injective nerve and the nerve of $\mathbf{Reach}_{G}$ are isomorphic and both contractible.
@@
If $\mathbf{C}=[n]$, then $N^{\iota}([n])=N([n])=\Delta^{n}$ is contractible. More generally, if $Q$ is any connected subgraph of $U([n])$, then its reachability $\mathbf{Reach}_{Q}$ is isomorphic to $[n]$, hence its (injective) nerve is also contractible. This case recovers the computation on transitive tournaments developed in Remark 3.8.
@@
Let $BK_{n}=\rho(K_{n})$ be the bicomplete directed graph on $n$ vertices. Then, its injective nerve is isomorphic to the (simplicial set associated to the) complex of injective words. More generally, if $G$ is any connected subgraph of $BK_{n}$ which is also a strongly connected component, then the reachability category $\mathbf{Reach}_{G}$ is such that $U(\mathbf{Reach}_{G})\cong BK_{n}$. Hence, $N^{\iota}(\mathbf{Reach}_{G})\simeq\Delta(V(BK_{n}))$ is a wedge of spheres. On the other hand, note that the usual nerve $N(\mathbf{Reach}_{G})$ is contractible.
Observe that we cannot expect the injective nerve of reachability categories to preserve products, not even up to homotopy:
@@
Consider the (Cartesian) product $\rho(K_{2})\times\rho(K_{2})$. Then, the associated injective nerve is isomorphic to the complex of injective words on four letters. Hence, up to homotopy, it is a wedge of $3$-spheres. On the other hand, we have $N^{\iota}(\rho(K_{2}))\simeq S^{1}$, and a Künneth theorem can not hold.
Likewise, we can not expect excision to generally hold, not even when the maps of reachability categories are Dwyer morphisms.
@@
Consider the spectral sequences associated to the filtration $F_{*}N(\mathbf{Reach}_{Q})\subseteq N(\mathbf{Reach}_{Q})$; this is called the magnitude-path spectral sequence [Asa23] (see also [DIMZ24] for the construction of the magnitude-path spectral sequence in the context of filtered simplicial sets). Likewise, we consider the spectral sequence associated to the filtration $F_{*}N^{\iota}(\mathbf{Reach}_{Q})\subseteq N^{\iota}(\mathbf{Reach}_{Q})$ of the injective nerve, which we call the regular magnitude-path spectral sequence. In the next result, we follow the notation of [Asa23].
@@
Let $G$ be a finite directed graph. The regular magnitude-path spectral sequence $IE_{*}^{*,*}$ associated to the filtration $F_{*}N^{\iota}(\mathbf{Reach}_{G})\subseteq N^{\iota}(\mathbf{Reach}_{G})$ satisfies the following properties:
@@
    The spectral sequence converges to the homology of the complex of injective words on $G$.
@@
Consider the filtration $F_{*}N^{\iota}(\mathbf{Reach}_{Q})$ of $N^{\iota}(\mathbf{Reach}_{Q})$. The 00-page of the spectral sequence $IE^{*,*}_{*}$ associated to this filtration is given by
$IE_{*}^{\ell,0}=\frac{F_{\ell}N^{\iota}(\mathbf{Reach}_{Q})}{F_{\ell-1}N^{\iota}(\mathbf{Reach}_{Q})}\ ,$
which is the eulerian chain complex. The induced differential $\partial\colon IE^{\ell,0}_{k}\to IE^{\ell,0}_{k-1}$ coincides with the differential of the eulerian magnitude chains, hence the isomorphism
$IE^{\ell,1}_{k}=\mathrm{EMH}_{k,\ell}(G)$
follows. Furthermore, when restricted to $k=\ell$, the induced second differential $\partial\colon IE^{\ell,1}_{k}\to IE^{\ell-1,1}_{k-1}$ computes the regular path homology groups of $G$ by [Asa23, Proposition 6.11] (adapted to the strongly regular elementary paths); that is, the homology of the chain complex $IE^{k-*,1}_{k-*}$ yields the regular path homology groups, proving the second item.
Observe that, by construction, when $G$ is finite the spectral sequence converges to the homology of $N^{\iota}(\mathbf{Reach}_{G})$. In view of Eq. (7), Lemma 5.2 and Proposition 4.6, the homology of $N^{\iota}(\mathbf{Reach}_{G})$ is the homology of the directed flag complex of $\mathrm{Pre}(G)$, that is the complex of injective words on $G$. To conclude, every morphism of filtrations yields a morphism of spectral sequences, which shows that the inclusions $F_{\ell}N^{\iota}(\mathbf{Reach}_{G})\subseteq F_{\ell}N(\mathbf{Reach}_{G})$ extend to a map of spectral sequences from the magnitude-path spectral sequence to $IE_{*}^{*,r}$. ∎
We now show some applications. First, as a consequence of Proposition 2.9 and of Theorem 5.13, we have that all graphs without 3- and 4-cycles have trivial path homology in degree $k\geq 2$:
@@
By Theorem 5.13, the regular path homology of $G$ is the homology of the chain complex $\mathrm{EMH}_{*,*}(G)$ appearing as diagonal in the second page of the regular magnitude-path spectral sequence. The result follows by Proposition 2.9. ∎
As a main consequence of Theorem 5.13, we provide computations of regular path homology groups of finite digraphs; in fact, the following holds:
@@
Let $G$ be a connected regularly diagonal finite directed graph. Then, the regular path homology of $G$ is the homology of $\mathrm{Inj}(G)=\mathrm{dFl}(\mathrm{Pre}(G))$.
@@
The regular magnitude-path spectral sequence converges to the homology of the complex of injective words $\mathrm{dFl}(\mathrm{Pre}(G))$. If $G$ is regularly diagonal, then the first page $IE^{1}_{*,*}$ is non-zero only in bidegrees $(\ell,k)$ with $k=\ell$, and similarly for the second page. Hence, the regular magnitude-path spectral sequence converges at the second page where all the non-trivial groups are concentrated on the diagonal. The result follows, as the diagonal coincides with the regular path homology of $G$. ∎

### v2

@@ 5. The injective nerve of categories and the regular magnitude-path spectral sequence
The goal of this section is to construct the regular magnitude-path spectral sequence, in analogy with the magnitude-path spectral sequence developed in [Asa23]. To do so, it shall be more convenient to use the categorical language, and to generalize the framework of eulerian magnitude homology to quivers. In the first subsection we introduce the tool of injective nerves of categories, and then we shall apply it to the reachable categories, getting the regular magnitude-path spectral sequence. As an application of the construction, we shall provide some computations of regular path homology of regularly diagonal graphs.
@@ 5.1. The injective nerve
Let $\mathbf{2}$ denote the category consisting of the objects $E$ and $V$, with two non-identity morphisms $s,t\colon E\to V$ called the source and the target. Let $\mathbf{Fin}$ be the full subcategory of $\mathbf{Set}$ of finite sets. By a (finite) quiver we shall mean a functor $Q\colon\mathbf{2}\to\mathbf{Fin}$. We shall also represent, and identify, a quiver $Q$ by the corresponding directed graph with set of vertices $V$, and set of edges $E$. Morphisms of quivers are natural transformations of functors. The category $\mathbf{Quiver}$ of finite quivers and morphisms of quivers is the functor category $\mathbf{Fun}(\mathbf{2},\mathbf{Fin})$.
@@
between the category of (finite) quivers and the 1-category $\mathbf{Cat}$ of small (finite) categories and functors. In this adjunction, the functor $\mathrm{Free}$ is the free functor that turns a quiver into a category, and $U$ is the corresponding forgetful functor.
Recall from Section 4 that an ordered $k$-clique of a directed graph $G$ is a totally ordered $k$-tuple $(v_{1},...,v_{k})$ of vertices of $G$ with the property that, for every $i<j$, the pair $(v_{i},v_{j})$ is an ordered edge of $G$. This definition extends to quivers in the natural way and, for any given quiver $Q$, we can construct the associated directed flag complex $\mathrm{dFl}(Q)$ as well. Recall also that the directed flag complex of a quiver is an ordered simplicial complex, and its simplices are determined by the vertices.
Let $\Delta\mathbf{Set}$ be the category of semi-simplicial sets, and $\mathbf{sSet}$ the category of simplicial sets. Recall that a semi-simplicial set is a sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of sets together with functions called face maps between them which encode that an element in $X_{n+1}$ has $n+1$ faces (i.e. boundary segments), and that such faces are elements of $X_{n}$; similarly, a simplicial set is a collection of sets $\{S_{n}\}_{n\in\mathbb{N}}$ together with face maps $f_{i}\colon S_{i+1}\to S_{i}$ and degeneracy maps $s_{i}\colon S_{i}\to S_{i+1}$, where the $f_{i}$’s are constructed similarly to the semi-simplicial case, and the $s_{i}$’s encode which $(i+1)$-simplices are really just $n$-simplices regarded as degenerate $(n+1)$-simplices.
Remark 5.1.
An ordered simplicial complex $K$ yields a semi-simplicial set, hence a simplicial set. This is given by adding to $K$ all possible degenerate simplices. More precisely, the $n$-simplices of the simplicial set $\iota(K)$ associated to the order simplicial complex $K$ consist of tuples $v_{i_{0}},\dots,v_{i_{n}}]$ with $v_{i_{j}}\leq v_{i_{j+1}}$ and such that the set of vertices $\{v_{i_{0}},\dots,v_{i_{n}}\}$ spans a simplex of $K$ – see, e.g. [Fri12, Example 3.3]. Therefore, for every simplex $v_{i_{0}},\dots,v_{i_{m}}]$ of $K$ we have in $\iota(K)$ all the (degenerate) simplices of the form
$v_{i_{0}},\dots,v_{i_{0}},\dots,v_{i_{1}},\dots,v_{i_{1}},\dots,v_{i_{m}},\dots,v_{i_{m}}]\ ,$
for any number of repetitions of each of the vertices. The non-degenerate simplices of $\iota(K)$ are in bijection with the simplices of $K$.
In view of Remark 5.1, for a small category $\mathbf{C}$, we can consider the following simplicial set:
(7) $N^{\iota}(\mathbf{C})\coloneqq\iota(\mathrm{dFl}(U(\mathbf{C})))\ ,$
where $U\colon\mathbf{Cat}\to\mathbf{Quiver}$ is the forgetful functor.
@@
For a small category $\mathbf{C}$, the simplicial set $N^{\iota}(\mathbf{C})$ is called the injective nerve of $\mathbf{C}$.
We can characterize the non-degenerate simplices of the injective nerve as the functors that are injective on objects. To see it, first consider the category $[n]$, that is the category with objects the natural numbers $0,\dots,n$ and with morphisms $i\to j$ if and only if $i\leq j$. Saying otherwise, $[n]$ is the free category generated by the transitive tournament $T_{n}$. Denote by $\mathbf{Cat}^{\iota}(\mathbf{D},\mathbf{C})$ the family of functors from $\mathbf{D}$ to $\mathbf{C}$ which are injective on objects.
@@
Recall the adjunction of Eq. (6). The set of ordered $(n+1)$-cliques of the quiver $U(\mathbf{C})$ is given by the ordered tuples $(c_{0},\dots,c_{n})$ consisting of objects of $\mathbf{C}$, with the further requirement that for all $i<j$ there is a morphism $f_{i,j}\colon c_{i}\to c_{j}$ in $\mathbf{C}$. Then, each such tuple defines a functor $S\colon[n]\to\mathbf{C}$ with $S(i)\coloneqq c_{i}$ and $S(i<j)\coloneqq f_{i,j}$. The functor $S$ is injective on objects because the set of vertices is ordered (and without repetitions). Vice versa, every functor $S\colon[n]\to\mathbf{C}$ which is injective on objects yields an ordered $(n+1)$-clique, hence an $n$-simplex of the directed flag complex – and, consequently, a non-degenerate simplex of $N^{\iota}(\mathbf{C})$. ∎
We now shall justify the term injective nerve used in Definition 5.2. Let $\Delta$ be the simplex category, and denote by $\Delta_{+}$ the wide subcategory of $\Delta$ containing only the injective functions. The category $\Delta_{+}$ is (equivalent to) the category of finite totally ordered sets and order-preserving injections. Recall that for a given (small) category $\mathbf{C}$, the nerve $N(\mathbf{C})$ of $\mathbf{C}$ is the simplicial set
@@
where $\mathbf{Cat}(-,\mathbf{C})$ denotes the Hom-functor. If we restrict to the category $\Delta_{+}$, we get the nerve of $\mathbf{C}$ as a semi-simplicial set – which means, without taking degeneracies. We shall further restrict it to the set of functors $\mathbf{Cat}^{\iota}(-,\mathbf{C})$ which are injective on objects; hence, we consider the composition
@@
The following result justifies the term of injective nerve used so far;
@@
The injective nerve $N^{\iota}(\mathbf{C})$ of a small category $\mathbf{C}$ is the simplicial set associated to
@@
By Lemma 5.3, the set of non-degenerate simplices of the injective nerve of a small category $\mathbf{C}$ can be described as the set of functors $n]\to\mathbf{C}$ which are injective on objects. The $i$-th face map operator in $\Delta_{+}$ acts by forgetting the $i$-th entry in the tuples $(c_{0},\dots,c_{k})$. Hence, the face maps are coherent with the face maps of the directed flag complex construction, yielding an isomorphism of semi-simplicial sets. The statement follows after considering the associated simplicial sets. ∎
As a consequence, Proposition 5.4 can be taken as alternative (categorical) definition of the injective nerve; in the following, we shall use both viewpoints.
@@
Recall that the nerve of a category is a fully faithful functor $N\colon\mathbf{Cat}\to\mathbf{sSet}$. In view of Proposition 5.4, also the injective nerve $N^{\iota}$ is a faithful functor. However, it is generally not full.
One may ask whether the injective nerve of a category is a (simplicial) nerve, at the cost of changing the base category. However, the following result tells us that in general the injective nerve is not the nerve of any category – nor of any quasicategory.
@@
Consider the category with two objects $0,1$, and non-trivial morphisms $0\to 1$, $1\to 0$. Consider the inner horn $\lambda^{n}_{i}$ in the case $n=2$, $i=1$ given by the arrows $0\to 1$, $1\to 0$. Then, $\lambda^{2}_{1}$ has no lift to $\Delta^{2}$ – because the simplex $(0,1,0)$ does not belong to $N^{\iota}(\mathbf{C})$. ∎
Observe that the inclusion $\mathbf{Cat}^{\iota}(-,\mathbf{C})\subseteq\mathbf{Cat}(-,\mathbf{C})$ induces a natural transformation
$N^{\iota}(\mathbf{C})\to N(\mathbf{C})\ .$
It would be interesting to know when $N^{\iota}(\mathbf{C})$ and $N(\mathbf{C})$ are homotopy equivalent. In complete analogy with Remark 4.8, we have the following partial result. Recall that each poset can be seen as a category in the standard way. Let $\mathbf{Poset}$ be the category of (finite) posets, and poset maps. There is an inclusion functor $\mathbf{Poset}\hookrightarrow\mathbf{Cat}$. The (injective) nerve of a poset $P$ is then the (injective) nerve of $P$, seen as a category.
Proposition 5.7.
Let $P$ be a poset. Then, the inclusion $\mathbf{Cat}^{\iota}(-,P)\subseteq\mathbf{Cat}(-,P)$ induces an isomorphism $N^{\iota}(P)\to N(P)$ of simplicial sets.
Proof.
The degenerate $n$-simplices of the nerve $N(P)$ consist of tuples of composable arrows in $P$, at least one of whose components is an identity morphism; the non-degenerate $n$-simplices are in bijection with the ordered tuples $x_{0},\dots,x_{n}]$ of elements of $P$, with $x_{i}<x_{i+1}$. Since $P$ is a poset, it does not contain loops, hence all objects appearing in a non-degenerate simplex $x_{0},\dots,x_{n}]$ are in fact distinct; i.e. $x_{i}\neq x_{j}$ for $i\neq j$. As a consequence, the set of non-degenerate $n$-simplices of the nerve $N(P)$ is in bijection with the set of functors $n]\to P$ which are injective on objects. In view of Lemma 5.3, this means that the inclusion $\mathbf{Cat}^{\iota}(-,P)\subseteq\mathbf{Cat}(-,P)$ induces a bijection between the non-degenerate simplices of $N(P)$ and of $N^{\iota}(P)$. As the face maps and degeneracies are defined in a coherent way, this is enough to prove that $N^{\iota}(P)\to N(P)$ is an isomorphism of simplicial sets. ∎
@@ 5.2. The regular magnitude-path spectral sequence
With en eye towards the regular magnitude-path spectral sequence, we are interested in the injective nerve of reachability categories, whose definition we now recall.
Let $Q$ be a directed graph or, more generally, a quiver.
@@
[…]
@@
Example 5.9.
If $Q$ is a directed tree, then $\mathrm{Free(Q)}$ and $\mathbf{Reach}_{Q}$ are isomorphic categories. More generally, $\mathrm{Free(Q)}$ and $\mathbf{Reach}_{Q}$ are isomorphic if and only if the quiver $Q$ does not contain directed cycles nor quasi-bigons – see [CR24, Proposition 3.8].
Using the description of the injective nerve of Remark 5.1, we can concretely describe the simplicial set $N^{\iota}(\mathbf{Reach}_{Q})$ as the simplicial set associated to the directed flag complex $\mathrm{dFl}(U(\mathbf{Reach}_{Q}))$. The directed flag complex of $U(\mathbf{Reach}_{Q})$ has the vertices of $Q$ as 0-simplices, and as $p$-simplices the (ordered) tuples $(x_{0},\dots,x_{p})$ of distinct vertices of $Q$ with the property that for each $i<j$ there is a directed path in $Q$ from $x_{i}$ to $x_{j}$. By transitivity, this means that the non-degenerate simplices of the injective nerve $N^{\iota}(\mathbf{Reach}_{Q})$ are given by tuples $(x_{0},\dots,x_{p})$ of distinct vertices of $Q$ with the property that for each $i=0,\dots,p-1$ there is a directed path in $Q$ from $x_{i}$ to $x_{i+1}$.
@@
Let $G$ be a directed graph. Then, from the definitions of injective nerve of a category (Remark 5.1 and Definition 5.2) and of reachability category (Definition 5.8) we get an isomorphism $U(\mathbf{Reach}_{G})\cong\mathrm{Pre}(G)$. As a consequence, by Proposition 4.7, we have
$N^{\iota}(\mathbf{Reach}_{G})=\iota(\mathrm{dFl}(U(\mathbf{Reach}_{G}))\cong\iota(\mathrm{dFl}(\mathrm{Pre}(G))\cong\iota(\mathrm{Inj}(G))\ ,$
and the injective nerve of $\mathbf{Reach}_{G}$ is isomorphic to the simplicial set associated to the complex of injective words on $G$.
Recall that, for a (ordered) simplicial complex $K$, the geometric realization $|\iota(K)|$ of the associated simplicial set returns the geometric realization of $K$.
@@
If $\mathbf{C}=[n]$, then $|N^{\iota}([n])|\cong|N([n])|$ is contractible (and homeomorphic to the standard simplex $\Delta^{n}$). More generally, if $Q$ is any connected subgraph of $U([n])$, then its reachability $\mathbf{Reach}_{Q}$ is isomorphic to $[n]$, hence its (injective) nerve is also contractible. This case recovers the computation on transitive tournaments developed in Remark 3.8.
@@
Let $BK_{n}=\rho(K_{n})$ be the bicomplete digraph on $n$ vertices. In view of Example 4.5, Remark 4.6, and Proposition 4.7, the injective nerve $N^{\iota}(\mathbf{Reach}_{BK_{n}})$ is isomorphic to the (simplicial set associated to the) classical complex of injective words. More generally, this holds if $G$ is any connected subgraph of $BK_{n}$, which is also strongly connected. Hence, $|N^{\iota}(\mathbf{Reach}_{G})|\simeq|\Delta(V(BK_{n}))|$ is homotopy equivalent to a wedge of spheres of dimension n1n-1 by Theorem 4.2. On the other hand, the nerve $N(\mathbf{Reach}_{G})$ is contractible.
In parallel with Remark 1.7, we can not expect the injective nerve of reachability categories to preserve products, not even up to homotopy:
@@
Consider the (Cartesian) product $\rho(K_{2})\times\rho(K_{2})$. Then, the injective nerve of the associated reachability category is isomorphic to the (simplicial set associated to the) complex of injective words on four letters. Hence, up to homotopy, it is a wedge of $3$-spheres. On the other hand, we have $|N^{\iota}(\mathbf{Reach}_{\rho(K_{2})})|\simeq S^{1}$.
For similar reasons as described in Example 5.13, we can not expect excision to generally hold, not even when the morphisms between reachability categories are Dwyer morphisms.
@@
Consider the spectral sequences associated to the filtration $F_{*}N(\mathbf{Reach}_{Q})\subseteq N(\mathbf{Reach}_{Q})$; this is called the magnitude-path spectral sequence [Asa23] (see also [DIMZ24] for the construction of the magnitude-path spectral sequence in the context of filtered simplicial sets). Likewise, we consider the spectral sequence associated to the filtration $F_{*}N^{\iota}(\mathbf{Reach}_{Q})\subseteq N^{\iota}(\mathbf{Reach}_{Q})$ of the injective nerve, which we call the regular magnitude-path spectral sequence. In the next result, we follow the notation of the pages of the spectral sequences as in [Asa23].
@@
Let $G$ be a directed graph. The regular magnitude-path spectral sequence $IE_{*}^{*,*}$ associated to the filtration $F_{*}N^{\iota}(\mathbf{Reach}_{G})\subseteq N^{\iota}(\mathbf{Reach}_{G})$ satisfies the following properties:
@@
    It converges to the homology of the complex of injective words on $G$.
@@
Consider the filtration $F_{*}N^{\iota}(\mathbf{Reach}_{G})$ of $N^{\iota}(\mathbf{Reach}_{G})$. For a simplicial set $S$, consider its normalized chain complex $\mathrm{N}(S)$. If $S$ is filtered, then also the normalized chain complex of $S$ is filtered, with filtration given by $F_{l}\mathrm{N}(S)=\mathrm{Im}(\mathrm{N}(F_{l}S)\to\mathrm{N}(S))$. The 0-page of the spectral sequence $IE^{*,*}_{*}$ associated to this filtration is given by the quotient
$IE_{*}^{\ell,0}=\frac{F_{\ell}\mathrm{N}(N^{\iota}(\mathbf{Reach}_{G}))}{F_{\ell-1}\mathrm{N}(N^{\iota}(\mathbf{Reach}_{G}))}$
of chain complexes – see also [DIMZ24]. By construction, $IE_{*}^{\ell,0}$ is isomorphic to the eulerian chain complex $\mathrm{EMC}_{*,\ell}(G)$. The induced differential $\partial\colon IE^{\ell,0}_{k}\to IE^{\ell,0}_{k-1}$ coincides with the differential of the eulerian magnitude chains, hence the isomorphism
$IE^{\ell,1}_{k}\cong\mathrm{EMH}_{k,\ell}(G)$
[…]
By construction, the considered spectral sequence converges to the homology of the injective nerve $N^{\iota}(\mathbf{Reach}_{G})$. In view of Remark 5.10, the homology of $N^{\iota}(\mathbf{Reach}_{G})$ is the homology of the complex of injective words on $G$.
To conclude, every morphism of filtrations yields a morphism of spectral sequences, which shows that the inclusions $F_{\ell}N^{\iota}(\mathbf{Reach}_{G})\subseteq F_{\ell}N(\mathbf{Reach}_{G})$ extend to a map of spectral sequences from the magnitude-path spectral sequence to $IE_{*}^{*,r}$. ∎
We now show some applications. First, observe that we can apply Theorem 5.15 to undirected graphs as well – by replacing an undirected graph $G$ with $\rho(G)$. As a consequence of Proposition 2.9 and of Theorem 5.15, we have that all graphs without 3- and 4-cycles have trivial regular path homology in degree $k\geq 2$:
@@
[…]
As a main consequence of Theorem 5.15, we provide computations of regular path homology groups of digraphs; in fact, the following holds:
@@
Let $G$ be a connected regularly diagonal directed graph. Then, the regular path homology groups of $G$ are the homology groups of $\mathrm{Inj}(G)$.
@@
The regular magnitude-path spectral sequence converges to the homology of the complex of injective words $\mathrm{Inj}(G)\cong\mathrm{dFl}(\mathrm{Pre}(G))$. If $G$ is regularly diagonal, then the first page $IE^{1}_{*,*}$ of the regular magnitude-path spectral sequence is non-zero only in bidegrees $(\ell,k)$ with $k=\ell$, and similarly for the second page. Hence, the regular magnitude-path spectral sequence converges at the second page where all the non-trivial groups are concentrated on the diagonal. The result follows, as the diagonal coincides with the regular path homology of $G$. ∎